# Goal Reaching with Eikonal-Constrained Hierarchical Quasimetric Reinforcement Learning

### v1

@@ 5 Eikonal-Constrained Hierarchical QRL (Eik-HiQRL)
To retain the advantages of Eik-QRL while mitigating its limitations, we propose a hierarchical architecture, yielding the Eik-HiQRL algorithm illustrated in Fig. 2. This design is motivated by the following considerations. First, as highlighted in previous work (park2024hiql), a key challenge in GCRL is the low signal-to-noise ratio in value estimates, especially in long-horizon tasks where the initial state and goal are far apart. park2024hiql propose a hierarchical solution in which a high-level actor outputs subgoals and a low-level controller attempts to reach them. Building on this, we show in Appendix B that quasimetric-parameterized value functions can have an even stronger effect than hierarchy alone on mitigating this signal-to-noise problem, further justifying the use of quasimetric value learning strategies such as Eik-QRL, HJB-QRL, and QRL together with hierarchy.
Given this result, a natural question is whether one could rely solely on a single quasimetric value function for both high- and low-level policies. Unfortunately, enforcing quasimetric structures in high-dimensional state spaces is itself challenging as the approximation error of quasimetric parameterizations grows exponentially with the state-space dimension [^1] This result can be readily derived by extending wang2022improved and combining it with standard results on covering numbers (e.g., shalev2014understanding).. A natural way to address this problem is to reduce the state dimensionality before applying quasimetric projection, either by projecting states into an embedding space or by selecting a subset of task-relevant state variables on which to define subgoals. If, in this process, we ensure that the resulting abstract space evolves under approximately isotropic dynamics, we can then leverage not only the benefits of the quasimetric structure but also the advantages of the Eikonal formulation (Remark 4.5).
Eik-HiQRL is designed precisely to reconcile these aspects: (1) quasimetric projection for the high-level value in a low-dimensional abstract space, (2) PDE-based regularization via the Eikonal formulation, and (3) a hierarchical structure that improves the signal-to-noise ratio. At the high level, we define a lower-dimensional abstract space $\bar{\mathcal{S}}$ (e.g., the positions of the agent or task-relevant objects) where the regularity assumptions of Eik-QRL hold and the quasimetric projection is tractable. This enables learning a high-level value function that improves subgoal generation, going beyond the single-value design in HIQL (park2024hiql). These subgoals then guide the low-level controller more effectively, further mitigating the signal-to-noise ratio issue in value function learning.
@@ 6 Experiments
Figure 3: Environments from OGbench (park2024ogbench) used in our experiments.
Our experiments are primarily conducted in the Offline GCRL setting, which offers a more controlled testbed than the online counterpart for evaluating value function learning and out-of-distribution generalization. In the offline case, the dataset remains fixed throughout training, making it possible to directly assess an algorithm’s ability to generalize beyond the training distribution. By contrast, in the online setting it is difficult to disentangle whether performance fluctuations arise from fortunate exploration or genuine improvements in generalization. To this end, we base our experiments on the tasks from OGbench (park2024ogbench), summarized in Fig. 3. We provide in Appendix E more details on these environments, including a discussion of how closely their dynamics satisfy the assumptions stated in Section 4.
@@
A few notes on implementations are in order. Eik-QRL and HJB-QRL are both implemented with a hierarchical actor, consistent with our analysis in Appendix B. Whereas QRL follows the original implementation of wang2023optimal as presented in Fig. 1. As mentioned, Appendix H.1 illustrates an additional comparison where QRL is combined with a hierarchical actor (denoted QRL:hi) in order to isolate the contribution of PDE constraints versus hierarchy. All the details for these variants are provided in Appendix D. Finally, Eik-HiQRL follows the design introduced in Section 5, where, for all the presented navigation tasks, the abstract space $\bar{\mathcal{S}}$ is defined as the agent’s coordinate space.
@@
This second set of experiments is divided into two parts. The first evaluates Eik-HiQRL against established Offline GCRL algorithms in the humanoidmaze navigation tasks. The second examines scenarios where the regularity assumptions do not hold, including MDPs with third-party objects and categorical variables in the state space. These experiments are designed to test performance in settings where our theoretical analysis is difficult to apply. The baselines considered in this section include Hierarchical Implicit Q-Learning (HIQL) (park2024hiql), Contrastive RL (CRL) (eysenbach2022contrastive), QRL (wang2023optimal), and Eikonal-regularized HIQL (Eik-HIQL) (giammarino2025physics). More details about these baselines, including how they differ from our algorithm, are provided in Appendix F.
@@ A.1 Assumptions restated
Lower bound. This part of the proof follows from Lemma 1 in wang2023optimal. In our case, however, assumptions $(A1)$ and $(A2)$ can only be satisfied simultaneously under the conditions stated in Lemma A.5.
@@ Appendix B Didactic example: joint benefits of hierarchy and quasimetric
In the following, we state Proposition B.1, which provides a novel bound on the probability of actor’s error under quasimetric value function parameterization. We provide a complete proof for Proposition B.1, and include a brief comparison with the corresponding results obtained without quasimetric projection in park2024hiql.
@@
The proof follows similarly to (park2024hiql, Proposition 4.1), however under the approximate quasimetric projection model.
@@
Without quasimetric projection from (park2024hiql, Proposition 4.1) we get:
@@ Appendix C Eik-HiQRL
Figure 7: Diagram of the Eik-HiQRL architecture. The high-level planning module consists of a quasimetric model $d^{h}$, parameterized as an IQE (wang2022improved), which is trained to optimize the physics-informed objective in (17), and a high-level policy $\pi^{h}$ trained via the advantage-weighted objective in (19). The low-level module comprises a value function $V^{l}$ and a goal representation network $\phi$, both instantiated as MLPs and trained to minimize the Implicit V-Learning loss in (18). The low-level policy $\pi^{l}$ is trained to maximize the advantage-weighted regression objective in (20).
@@ C.1 Overview
Refer to Fig. 7 for a graphical illustration of the Eik-HiQRL architecture. The high-level planning module consists of a quasimetric model parameterized by an IQE (wang2022improved), together with a high-level policy that outputs a subgoal embedding $z_{t+k}=\phi([s_{t+k},s_{t}])$. The low-level control module includes a value function $V^{l}$ parameterized by an MLP, as well as a goal representation network $\phi$, also instantiated as an MLP with length normalization applied to its final layer. The low-level policy takes the current state and the subgoal embedding as input and outputs an action to interact with the environment.
@@
where $z_{\phi}$ and $z^{\prime}_{\phi}$ are the goal embeddings produced by $\phi_{\bm{\theta}_{\phi}}$, $\bar{\bm{\theta}}_{V}$ are the parameters of the target value network (mnih2013playing), and $L_{2}^{\iota}(x)=|\iota-\mathbbm{1}(x<0)|x^{2}$ is the expectile loss with $\iota\in[0.5,1]$.
@@ Appendix E Offline GCRL Environments used in Section 6
Figure 9: Environments from OGbench (park2024ogbench) used in our experiments in Section 6.
@@ Appendix F Offline GCRL Baselines used in Section 6
Table 4 summarizes the main structural differences among Eik-QRL (8), QRL (wang2023optimal), HIQL (park2024hiql), Eik-HIQL (giammarino2025physics), CRL (eysenbach2022contrastive) and Eik-HiQRL (Fig. 2).
QRL learns a quasimetric goal-conditioned value function without any PDE-based constraints. Eik-QRL builds on QRL by adding Eikonal constraints, which provides similar theoretical guarantees under our regularity assumptions and a smoother value landscape, while also introducing a hierarchical actor. HIQL, in contrast, uses a hierarchical actor but learns a standard TD-based value (rather than a quasimetric) and does not impose any PDE constraints. Eik-HIQL augments HIQL with Eikonal regularization, yet it maintains a non-quasimetric value and does not decompose the value hierarchically. Our Eik-HiQRL combines the strengths of these approaches: it uses a quasimetric value constrained by the Eikonal PDE at the high level, together with both a hierarchical actor and a hierarchical value structure. This design preserves the theoretical benefits of Eik-QRL where its assumptions are most plausible, while mitigating its practical limitations through hierarchy and enabling improved scaling to long-horizon tasks. Finally, CRL (eysenbach2022contrastive) is a substantially different method compared to the previous algorithms which learns a critic based on contrastive learning principles. We include it here for the sake of completeness.
@@ H.3 Full Comparison Eik-HiQRL vs Baselines
Table 7: Summary of our experimental results. All agents are trained for $10^{6}$ steps using 10 seeds and evaluated every $10^{5}$ steps. We report the mean and standard deviation across seeds for the best evaluation. For each seed, evaluations are conducted over 55 different random goals, as designed in park2024ogbench, with the learned policy tested for $50$ episodes per goal. Results within $95\%$ of the best value are written in bold.
@@ H.5 Trajectory-free Experiments
To investigate the trajectory-free regime, we evaluate Eik-QRL on a robotic goal-reaching task based on the FetchReach environment (1802.09464). The setup, illustrated in Fig. 27(b), uses a 7-DoF Fetch mobile manipulator equipped with a two-finger parallel gripper; the underlying environment controls the robot in Cartesian space by specifying small displacements of the end-effector, while inverse kinematics are handled internally by the simulator. To make this setting fully trajectory-free, we restrict the state space to the end-effector Cartesian coordinates $(x,y,z)$, and randomly sample $100\text{k}$ end-effector positions from the robot workspace, aggregating them into a replay buffer without storing any transition trajectories.
@@ H.6.1 Pointmaze
This Pointmaze environment consists of a 2-DoF point mass (a “ball”) that is force-actuated in the Cartesian directions $x$ and $y$, and must reach a target goal within a closed maze (fu2020d4rl). The environment can be instantiated with different maze layouts and increasing levels of difficulty, adapted from fu2020d4rl. We consider multiple maze sizes and configurations, including single-goal and multi-goal variants, where the latter are referred to as diverse mazes. A further group of variations instantiates another class of diverse mazes in which both goal locations and initial agent states are randomly sampled at each reset. These environments share the same base identifiers as their default counterparts, with the suffix Diverse_GR (GR stands for Goal and Reset) appended.
@@ H.6.2 FetchReach
We further evaluate QRL and Eik-QRL in an online GCRL setting on the FetchReach environment introduced in 1802.09464. The task is to move the end effector of a 7-DoF Fetch manipulator, equipped with a two-finger parallel gripper, to a randomly sampled target position within the robot workspace. The robot is controlled in Cartesian space: each action specifies a small displacement $(\Delta x,\Delta y,\Delta z)$ of the end effector together with a scalar command that opens or closes the gripper, while inverse kinematics are handled internally by the simulator. The resulting continuous dynamics are smooth and approximately isotropic in the control space, so both Assumption 4.2 (Lipschitz dynamics) and, consequently, Assumption 4.4 (Lipschitz goal-conditioned value function) are expected to hold reasonably well in this setting.

### v2

@@ 5 Eikonal-Constrained Hierarchical QRL (Eik-HiQRL)
To retain the advantages of Eik-QRL while mitigating its limitations, we propose a hierarchical architecture, yielding the Eik-HiQRL algorithm illustrated in Fig. 2. This design is motivated by the following considerations. First, as highlighted in previous work (Park et al., 2024b), a key challenge in GCRL is the low signal-to-noise ratio in value estimates, especially in long-horizon tasks where the initial state and goal are far apart. Park et al. (2024b) propose a hierarchical solution in which a high-level actor outputs subgoals and a low-level controller attempts to reach them. Building on this, we show in Appendix B that quasimetric-parameterized value functions can have an even stronger effect than hierarchy alone on mitigating this signal-to-noise problem, further justifying the use of quasimetric value learning strategies such as Eik-QRL, HJB-QRL, and QRL together with hierarchy.
Given this result, a natural question is whether one could rely solely on a single quasimetric value function for both high- and low-level policies. Unfortunately, enforcing quasimetric structures in high-dimensional state spaces is itself challenging as the approximation error of quasimetric parameterizations grows exponentially with the state-space dimension [^1] This result can be readily derived by extending Wang and Isola (, Theorem 3) and combining it with standard results on covering numbers (e.g., Shalev-Shwartz and Ben-David (2014, Ex. 27.1)).. A natural way to address this problem is to reduce the state dimensionality before applying quasimetric projection, either by projecting states into an embedding space or by selecting a subset of task-relevant state variables on which to define subgoals. If, in this process, we ensure that the resulting abstract space evolves under approximately isotropic dynamics, we can then leverage not only the benefits of the quasimetric structure but also the advantages of the Eikonal formulation (Remark 4.5).
Eik-HiQRL is designed precisely to reconcile these aspects: (1) quasimetric projection for the high-level value in a low-dimensional abstract space, (2) PDE-based regularization via the Eikonal formulation, and (3) a hierarchical structure that improves the signal-to-noise ratio. At the high level, we define a lower-dimensional abstract space $\bar{\mathcal{S}}$ (e.g., the positions of the agent or task-relevant objects) where the regularity assumptions of Eik-QRL hold and the quasimetric projection is tractable. This enables learning a high-level value function that improves subgoal generation, going beyond the single-value design in HIQL (Park et al., 2024b). These subgoals then guide the low-level controller more effectively, further mitigating the signal-to-noise ratio issue in value function learning.
@@ 6 Experiments
Figure 3: Environments from OGbench (Park et al., 2024a) used in our experiments.
Our experiments are primarily conducted in the Offline GCRL setting, which offers a more controlled testbed than the online counterpart for evaluating value function learning and out-of-distribution generalization. In the offline case, the dataset remains fixed throughout training, making it possible to directly assess an algorithm’s ability to generalize beyond the training distribution. By contrast, in the online setting it is difficult to disentangle whether performance fluctuations arise from fortunate exploration or genuine improvements in generalization. To this end, we base our experiments on the tasks from OGbench (Park et al., 2024a), summarized in Fig. 3. We provide in Appendix E more details on these environments, including a discussion of how closely their dynamics satisfy the assumptions stated in Section 4.
@@
A few notes on implementations are in order. Eik-QRL and HJB-QRL are both implemented with a hierarchical actor, consistent with our analysis in Appendix B. Whereas QRL follows the original implementation of Wang et al. (2023) as presented in Fig. 1. As mentioned, Appendix H.1 illustrates an additional comparison where QRL is combined with a hierarchical actor (denoted QRL:hi) in order to isolate the contribution of PDE constraints versus hierarchy. All the details for these variants are provided in Appendix D. Finally, Eik-HiQRL follows the design introduced in Section 5, where, for all the presented navigation tasks, the abstract space $\bar{\mathcal{S}}$ is defined as the agent’s coordinate space.
@@
This second set of experiments is divided into two parts. The first evaluates Eik-HiQRL against established Offline GCRL algorithms in the humanoidmaze navigation tasks. The second examines scenarios where the regularity assumptions do not hold, including MDPs with third-party objects and categorical variables in the state space. These experiments are designed to test performance in settings where our theoretical analysis is difficult to apply. The baselines considered in this section include Hierarchical Implicit Q-Learning (HIQL) (Park et al., 2024b), Contrastive RL (CRL) (Eysenbach et al., 2022), QRL (Wang et al., 2023), and Eikonal-regularized HIQL (Eik-HIQL) (Giammarino et al., 2025). More details about these baselines, including how they differ from our algorithm, are provided in Appendix F.
@@ A.1 Assumptions restated
Lower bound. This part of the proof follows from Lemma 1 in Wang et al. (2023). In our case, however, assumptions $(A1)$ and $(A2)$ can only be satisfied simultaneously under the conditions stated in Lemma A.5.
@@ Appendix B Didactic example: joint benefits of hierarchy and quasimetric
In the following, we state Proposition B.1, which provides a novel bound on the probability of actor’s error under quasimetric value function parameterization. We provide a complete proof for Proposition B.1, and include a brief comparison with the corresponding results obtained without quasimetric projection in Park et al. (2024b).
@@
The proof follows similarly to (Park et al., 2024b, Proposition 4.1), however under the approximate quasimetric projection model.
@@
Without quasimetric projection from (Park et al., 2024b, Proposition 4.1) we get:
@@ Appendix C Eik-HiQRL
Figure 7: Diagram of the Eik-HiQRL architecture. The high-level planning module consists of a quasimetric model $d^{h}$, parameterized as an IQE (Wang and Isola, ), which is trained to optimize the physics-informed objective in (17), and a high-level policy $\pi^{h}$ trained via the advantage-weighted objective in (19). The low-level module comprises a value function $V^{l}$ and a goal representation network $\phi$, both instantiated as MLPs and trained to minimize the Implicit V-Learning loss in (18). The low-level policy $\pi^{l}$ is trained to maximize the advantage-weighted regression objective in (20).
@@ C.1 Overview
Refer to Fig. 7 for a graphical illustration of the Eik-HiQRL architecture. The high-level planning module consists of a quasimetric model parameterized by an IQE (Wang and Isola, ), together with a high-level policy that outputs a subgoal embedding $z_{t+k}=\phi([s_{t+k},s_{t}])$. The low-level control module includes a value function $V^{l}$ parameterized by an MLP, as well as a goal representation network $\phi$, also instantiated as an MLP with length normalization applied to its final layer. The low-level policy takes the current state and the subgoal embedding as input and outputs an action to interact with the environment.
@@
where $z_{\phi}$ and $z^{\prime}_{\phi}$ are the goal embeddings produced by $\phi_{\bm{\theta}_{\phi}}$, $\bar{\bm{\theta}}_{V}$ are the parameters of the target value network (Mnih, 2013), and $L_{2}^{\iota}(x)=|\iota-\mathbbm{1}(x<0)|x^{2}$ is the expectile loss with $\iota\in[0.5,1]$.
@@ Appendix E Offline GCRL Environments used in Section 6
Figure 9: Environments from OGbench (Park et al., 2024a) used in our experiments in Section 6.
@@ Appendix F Offline GCRL Baselines used in Section 6
Table 4 summarizes the main structural differences among Eik-QRL (8), QRL (Wang et al., 2023), HIQL (Park et al., 2024b), Eik-HIQL (Giammarino et al., 2025), CRL (Eysenbach et al., 2022) and Eik-HiQRL (Fig. 2).
QRL learns a quasimetric goal-conditioned value function without any PDE-based constraints. Eik-QRL builds on QRL by adding Eikonal constraints, which provides similar theoretical guarantees under our regularity assumptions and a smoother value landscape, while also introducing a hierarchical actor. HIQL, in contrast, uses a hierarchical actor but learns a standard TD-based value (rather than a quasimetric) and does not impose any PDE constraints. Eik-HIQL augments HIQL with Eikonal regularization, yet it maintains a non-quasimetric value and does not decompose the value hierarchically. Our Eik-HiQRL combines the strengths of these approaches: it uses a quasimetric value constrained by the Eikonal PDE at the high level, together with both a hierarchical actor and a hierarchical value structure. This design preserves the theoretical benefits of Eik-QRL where its assumptions are most plausible, while mitigating its practical limitations through hierarchy and enabling improved scaling to long-horizon tasks. Finally, CRL (Eysenbach et al., 2022) is a substantially different method compared to the previous algorithms which learns a critic based on contrastive learning principles. We include it here for the sake of completeness.
@@ H.3 Full Comparison Eik-HiQRL vs Baselines
Table 7: Summary of our experimental results. All agents are trained for $10^{6}$ steps using 10 seeds and evaluated every $10^{5}$ steps. We report the mean and standard deviation across seeds for the best evaluation. For each seed, evaluations are conducted over 55 different random goals, as designed in Park et al. (2024a), with the learned policy tested for $50$ episodes per goal. Results within $95\%$ of the best value are written in bold.
@@ H.5 Trajectory-free Experiments
To investigate the trajectory-free regime, we evaluate Eik-QRL on a robotic goal-reaching task based on the FetchReach environment (Plappert et al., 2018). The setup, illustrated in Fig. 27(b), uses a 7-DoF Fetch mobile manipulator equipped with a two-finger parallel gripper; the underlying environment controls the robot in Cartesian space by specifying small displacements of the end-effector, while inverse kinematics are handled internally by the simulator. To make this setting fully trajectory-free, we restrict the state space to the end-effector Cartesian coordinates $(x,y,z)$, and randomly sample $100\text{k}$ end-effector positions from the robot workspace, aggregating them into a replay buffer without storing any transition trajectories.
@@ H.6.1 Pointmaze
This Pointmaze environment consists of a 2-DoF point mass (a “ball”) that is force-actuated in the Cartesian directions $x$ and $y$, and must reach a target goal within a closed maze (Fu et al., 2020). The environment can be instantiated with different maze layouts and increasing levels of difficulty, adapted from Fu et al. (2020). We consider multiple maze sizes and configurations, including single-goal and multi-goal variants, where the latter are referred to as diverse mazes. A further group of variations instantiates another class of diverse mazes in which both goal locations and initial agent states are randomly sampled at each reset. These environments share the same base identifiers as their default counterparts, with the suffix Diverse_GR (GR stands for Goal and Reset) appended.
@@ H.6.2 FetchReach
We further evaluate QRL and Eik-QRL in an online GCRL setting on the FetchReach environment introduced in Plappert et al. (2018). The task is to move the end effector of a 7-DoF Fetch manipulator, equipped with a two-finger parallel gripper, to a randomly sampled target position within the robot workspace. The robot is controlled in Cartesian space: each action specifies a small displacement $(\Delta x,\Delta y,\Delta z)$ of the end effector together with a scalar command that opens or closes the gripper, while inverse kinematics are handled internally by the simulator. The resulting continuous dynamics are smooth and approximately isotropic in the control space, so both Assumption 4.2 (Lipschitz dynamics) and, consequently, Assumption 4.4 (Lipschitz goal-conditioned value function) are expected to hold reasonably well in this setting.